Check whether or not S = { a0 + a1x + .... + an(x^n) belongs to Z[x] | 5 | a0} is an ideal of Z[x].
1
Expert's answer
2020-05-13T20:08:16-0400

Let "f\in\mathbb Z[x],\,\,f=b_0+b_1x+\dots" and "s\in S"

"fs=a_0b_0+\dots \in \mathbb Z[x]" and "5\mid a_0'=a_0b_0"

Hence, "fs\in S" and "S" is an ideal of "\mathbb Z[x]"